# Wholesale Market Participation of DERA: Competitive DER Aggregation

### v4

@@ III-B Properties of DERA Competitive with NEM X
${\cal K}_{n}=\zeta S_{n}^{\mbox{\tiny NEM}}(g_{n},\underline{C}_{n},\overline{C}_{n}),$ (12)
@@
which is used in (5) with profit ratio $\zeta\geq 1$ to obtain competitive aggregation over the DSO’s NEM-based aggregation with the same network access. [^7] Customers owning DERs switch from NEM X to DERA for higher consumer surplus, granting DERs control to DERA upon joining. In this subsection, the network access limits carry the subscript $n$, which is equivalent to $m$ since we set ${\cal N}_{m}=1$ and $K=1$ for simplicity.
@@
The proof is provided in Appendix VIII-C. Such price stability comes directly from the ${\cal K}$-competitive constraint, which enforces a lower bound for customer surplus and thus naturally limits the maximum customer payment. Note that the two-part pricing of Griddy [7] is not a stable pricing mechanism because the retail rate is tied directly to the real-time LMP.
@@
The proof is provided in Appendix VIII-D. If NEM is not crediting BTM DG of prosumer at a higher production rate $\pi^{-}$ than the expected wholesale LMP $\pi$, our aggregation method is profitable while competing with NEM to attain customers. In practice, $\pi^{-}\leq\mathbb{E}[\pi]\leq\pi^{+}$ happens since many states including California are setting the export rate $\pi^{-}$ at the avoided cost rate, which is, in many cases close to the expected LMP $\pi$, to mitigate cross-subsidies [33, 6, 4].
@@ IV-A Offer/Bid Curves of DERA in Energy Markets
Let $Q$ be DERA’s aggregated quantity to buy (when $Q<0$) or sell (when $Q>0$) and $\pi$ be the wholesale market LMP. Let $G:=\sum_{n=1}^{N}g_{n}$ be the BTM DG aggregated by DERA. In a competitive market, a price-taking DERA participant bids truthfully with its aggregated supply function
@@
where ${\bm{d}}^{*}_{n}$ is defined in (7). Note that the inverse of the DERA supply function $\mathscr{F}^{-1}(Q)$ defines the offer/bid curves of the DERA. For a quantity bid, the DERA forecasts the LMP and computes the optimal net production with (13). In contrast, for a price-quantity bid/offer curve $\mathscr{F}^{-1}(Q)$, the DERA avoids LMP forecasting, as the ISO clears the market using the submitted curve and ensures consistency between the LMP and the resulting dispatch. This is further explained in Lemma 1 below. Note also that the supply function depends on the aggregated BTM generation $G$, which is not known to the DERA at the time of the market auction. In practice, $G$ can be approximated by using historical data or $N\mathbb{E}(g_{n})$ via the Law of Large Numbers involving $N$ independent prosumers or via the Central Limit Theorem for independent and dependent random variables [34].
@@ IV-B Market Efficiency with DERA Participation
Consider a transmission network with $I$ buses. At each bus of the transmission network, we assume $M$ PoAs at the distribution network are connected and $N$ prosumers are aggregated by the proposed DERA model. Denote $U_{in}$ as the concave utility function for the $n$-th prosumer at the $i$-th transmission network bus. ${\bm{g}}_{n}:=(g_{in})_{i\in[I]}$ and ${\bm{d}}_{n}:=(d_{in})_{i\in[I]}$ are respectively the vectors of BTM DG generation and energy consumption for the prosumers. For simplicity, we ignore the number of energy-consuming devices for each prosumer, i.e., $K=1$, in this section. At each transmission bus, we sum up all load-serving entities and generators into one demand function and supply function. The load-serving entity at bus $i$ purchases electricity $e_{i}$ with a concave benefit function $B_{i}(e_{i})$. The generator at bus $i$ produces $p_{i}$ with a convex cost function $C_{i}(p_{i})$. Denote ${\bm{e}}:=(e_{i})_{i\in[I]}$, ${\bm{p}}:=(p_{i})_{i\in[I]}$. ${\bm{f}}\in\mathbb{R}^{L}$ is the line flow limit for $L$ branches of the transmission network. ${\bm{S}}\in\mathbb{R}^{L\times I}$ is the network parameter for DC power flow model.
@@
When prosumers participate in the wholesale market indirectly through the proposed DERA model with offer/bid curve (13), social welfare $\mbox{\sf SW}_{\mbox{\rm\tiny DERA}}$ is the optimal value of convex problem
@@
[…]
@@
where ${\cal D}_{in}:=[\max\{\underline{d}_{in},g_{in}-\overline{c}_{in}\},\min\{\bar{d}_{in},g_{in}+\underline{c}_{in}\}]$. The access limits $\overline{c}_{in}$ and $\underline{c}_{in}$ represent the distribution network injection and withdrawal capacities allocated to each prosumer. [^8] Detail formulations are in (51)(52) of the appendix. These values are consistent with those computed in (7), ensuring a fair comparison. Under this setup, prosumers participating directly in the wholesale market face the same access constraints at each PoA as those in our proposed DERA model. By solving (16), we obtain the bid/offer curve for prosumer $n$ at bus $i$, i.e.,
@@
The sketch of proof is provided in Appendix VIII-F, which relies on the fact that the proposed DERA has its bidding curve (13) equal to the sum of the prosumer’s bidding curve in (17). From this, we can establish that the wholesale market clearing problem with the direct participation of all prosumers has the same market-clearing results as (14).
@@ V DERA-DSO coordination
All generation and consumption resources aggregated by DERA need to bypass the distribution network to participate in the wholesale market. The DERA aggregation presented in this work ensures that the aggregated DER at each distribution network PoA is bounded by access limits imposed through the distribution network access limit auction in [25]. A DERA submits a bid curve in this auction representing its willingness to acquire access at PoAs in the distribution system.
We assume that a DERA is a price taker in the access limit auction. Therefore, the bid-in demand curve for network access from the DERA at a particular PoA is the marginal benefit (profit) from having a DER aggregation under the PoA. The maximum expected profit of DERA is given by
@@ V-A DERA Benefit Function for Distribution Network Access
Denote ${\bm{h}}_{n}(x):=\sum_{k=1}^{K}h_{nk}(x)$ and two thresholds
We have the following form for DERA’s benefit function.
@@
The sketch of proof is provided in Appendix VIII-A. Here, we have $U_{n}({\bm{h}}_{n}(x))=\sum_{k=1}^{K}U_{nk}(h_{nk}(x))$ from the additive property of the utility function. The optimal DERA surplus is decomposed into three terms: the surplus dependent on the withdraw access $\underline{\phi}_{m}({\underline{C}_{m}})$, the surplus dependent on the injection access $\overline{\phi}_{m}({\overline{C}})$, and the surplus independent of the network access. $\varphi({\underline{{\bm{C}}}},{\overline{{\bm{C}}}})$ is separable across injection and withdrawal access over $M$ PoAs. Therefore, DERA can bid separately for the distribution network accesses at different PoAs when coordinating with DSO. At PoA $m$ with less BTM DG, i.e., $\sum_{n=1}^{{\cal N}_{m}}g_{n}\leq\underline{q}_{m}$, DERA’s benefit depends on the withdrawal access. Conversely, if there is more BTM DG, i.e., $\overline{q}_{m}\leq\sum_{n=1}^{{\cal N}_{m}}g_{n}$, DERA’s benefit depends on the injection access. Related simulations are in Sec. VI-D.
@@ V-B Long-Run Equilibrium for Competitive DERA
In a long-run competitive industry, we explore how many DERA can survive. DERAs compete to attract customers, attain distribution network access, and participate in the wholesale market. We assume all DERAs adopt the competitive DER aggregation method in (6). The condition for a competitive long-run equilibrium [35, P193] has two components: (i) the marginal benefit of DERA equals the marginal cost of DSO for providing the distribution network access, and (ii) all DERAs have profits equal to zero, i.e., DERA’s profit in conducting aggregation equals DERA’s payment to acquire distribution network access. Related derivations and simulations are in Sec. VI-E and Appendix VIII-G.
@@ VI-A Parameter Settings
where $\alpha=\$0.4/\mbox{kWh},\beta=\$0.1/(\mbox{kWh})^{2}$ [27]. Let the marginal utility $V^{-1}\in[\underline{d},\bar{d}]$ for the consumption limits. [^9] We simulated the case with prosumer device number $K=1$ for simplicity.
We used NEMa and NEMp to represent the DER aggregation under NEM when prosumers were active and passive, respectively. Passive customers are not responsive to the retail prices, but active customers will optimize their energy consumption given the retail price and the BTM DG generations. Based on PG&E residential rate, we set $\pi^{+}=\$0.3/\mbox{kWh}$ for the NEM. We assumed $\pi^{-}=\mathbb{E}[\pi]$ and the fixed cost of NEM was covered by extracting fixed payment from DERA, so we simulated with $\pi^{0}=\$0$. Gao-Alshehri-Birge (GAB) represented the two-part pricing in [12], which allowed customers to sell BTM DG to the DERA while purchasing energy from its incumbent utility company. Detailed models for NEMa, NEMp, and GAB are provided in Appendix VIII-B. Our DER aggregation method was simulated in Co.NEMa and Co.GAB, competitive with NEMa and GAB, respectively. For Co.NEMa, we set profit ratio $\zeta$ at the upper bound in the proof of Proposition 2. For Co.GAB, we set $\zeta=1.05$ to provide 5% more customer surplus than the GAB competitor.
@@ VI-D Benefit Function of DERA for Distribution Network Access
Figure 5: DERA Benefit function $\varphi$. (Left: withdrawal access $-\underline{C}$; right: injection access $\overline{C}$.)
@@ VI-E Long-Run Competitive Equilibrium of DERAs
In the long-run competitive equilibrium analysis with multi-interval aggregation of DERs, we assumed 200 DERAs initially existed and computed the expected number of surviving DERAs in the long run. For simplicity, we assumed DERAs were homogeneous and had the same setting as Sec. VI-D. Prosumers had the same expected DG generation created from the 24-hour roof-top solar data in Pecan Street [37]. [^11] Detail DG trajectories and the long-run equilibrium results for single-interval aggregation are shown in Appendix VIII-G, providing intuitions about long-run equilibrium for multi-interval aggregation here. We multiplied the mean of 24-hour DG by ϵ1+subscriptitalic-$\epsilon_{1}\in\mathbb{R}_{+}$ to simulate different DG installation capacities and sampled 10,000 random DG scenarios. DERA submitted the benefit function, as in Fig. 5, to acquire hourly distribution network access. Same as [25], the DSO cost function for providing distribution network access was assumed to be the sum of quadratics, $J(x)=\frac{1}{2}bx^{2}+ax$ with $a=\$0.009/\mbox{kWh},b=\$0.0005/(\mbox{kWh})^{2}$ for both the injection and withdrawal access. We multipied DSO’s cost $J$ by ϵ2+subscriptitalic-$\epsilon_{2}\in\mathbb{R}_{+}$ to simulate different levels of DSO’s costs.
@@ VII Conclusions
A major challenge to realizing the goals of FERC Order $2$ is enabling profit-maximizing DERAs to effectively compete with the retail programs offered by distribution utilities. To this end, this paper considers the competitive DER aggregation of a profit-seeking DERA in the wholesale electricity market. As a wholesale market participant, DERA can both inject and withdraw power from the wholesale market. It is shown that the proposed DERA model maximizes its profit while providing competitive services to its customers with higher surpluses than those offered by the distribution utilities. We also establish that the resulting social welfare from DERA’s participation on behalf of its prosumers is the same as that gained by the direct participation of price-taking prosumers, making the proposed DERA aggregation model optimal in achieving wholesale market efficiency. Additionally, we derive two significant optimal price-quantity bids of DERA, of which one is submitted to the wholesale market, and the other to the distribution network access allocation mechanism [25].
An open issue of the proposed aggregation solution is that the payment functions for prosumers are nonlinear and non-uniform. Although each customer is guaranteed to be better off than the competing scheme, two customers producing the same amount may be paid and compensated differently. In other words, the total charge/credits depend not only on the quantity but also on the flexibilities of the demand and constraints imposed by the prosumer. Note that a profit-seeking DERA participating in the wholesale electricity market is not subject to the same regulation as a regulated utility. Such non-uniform pricing is acceptable and has also been proposed in the form of non-uniform fixed charges [12, 6].

### v5

@@ III-B NEM benchmarks
In summary, the prosumer surplus under NEM, $S_{\mbox{\tiny NEM}}(g,\underline{C},\overline{C})$ is given by
$S_{\mbox{\tiny NEM}}(g,\underline{C},\overline{C})=\begin{cases}S_{\mbox{\tiny NEM-a}}(g,\underline{C},\overline{C}),&\text{active prosumer},\\ S_{\mbox{\tiny NEM-p}}(g,\underline{C},\overline{C}),&\text{passive prosumer}.\end{cases}$ (16)
III-C Properties of DERA Competitive with NEM
We analyze the profitability of DERA and the energy consumption cost of aggregated prosumers when DERA is competitive with a regulated NEM tariff parameterized. Assume $0\leq\pi^{-}\leq\pi^{+}$ and customers’ surpluses under NEM are nonnegative [5]. From (47), we have the $n$-th prosumer surplus under NEM $S_{n}^{\mbox{\tiny NEM}}(g_{n},\underline{C}_{n},\overline{C}_{n})$, whose computation depends on the DG generation and network access limits. DERA sets the benchmark prosumer surplus
${\cal K}_{n}=\zeta S_{n}^{\mbox{\tiny NEM}}(g_{n},\underline{C}_{n},\overline{C}_{n}),~~\zeta\geq 1,$ (17)
@@ III-C Properties of DERA Competitive with NEM
which is used in (5) with profit ratio $\zeta$ to obtain competitive aggregation over the DSO’s NEM-based aggregation with the same network access. [^4] Customers owning DERs switch from NEM to DERA for higher consumer surplus, granting DERs control to DERA upon joining. In this subsection, the network access limits carry the subscript $n$, which is equivalent to $m$ since we set ${\cal N}_{m}=1$ and $K=1$ for simplicity.
@@
See proof in the appendix. Such price stability comes directly from the $\cal K$}-competitive constraint, which enforces a lower bound for customer surplus and thus naturally limits the maximum customer payment. Note that the two-part pricing of Griddy [8] is not a stable pricing mechanism because the retail rate is tied directly to real-time LMP.
@@
The proof is provided in the appendix. In practice, the condition $\pi^{-}\leq\mathbb{E}[\pi]\leq\pi^{+}$ is often satisfied. For instance, in many states—including California—the export rate $\pi^{-}$ is set near the avoided cost, which typically approximates the expected LMP $\mathbb{E}[\pi]$, as a way to mitigate cross-subsidies [7].
@@ IV-A Offer/Bid Curves of DERA in Energy Markets
Let $Q$ be DERA’s aggregated quantity to buy (when $Q<0$) or sell (when $Q>0$), and $\pi$ be the wholesale market LMP. Let $G:=\sum_{n=1}^{N}g_{n}$ be the BTM DG aggregated by DERA. In a competitive market, a price-taking DERA participant bids truthfully with its aggregated supply function
@@
where ${\bm{d}}^{*}_{n}$ is defined in (10). Note that the inverse of the DERA supply function $\mathscr{F}^{-1}(Q)$ defines the offer/bid curves of the DERA. For a quantity bid, the DERA forecasts the LMP and computes the optimal net production with (18). In contrast, for a price-quantity bid/offer curve $\mathscr{F}^{-1}(Q)$, the DERA avoids LMP forecasting, as the ISO clears the market using the submitted curve and ensures consistency between the LMP and the resulting dispatch. More details are provided in Lemma 1, and a simulation of this offer/bid curve is presented in our previous paper [1].
Note also that the supply function depends on the aggregated BTM generation $G$, which is not known to the DERA at the time of the market auction. In practice, $G$ can be approximated by using historical data or $N\mathbb{E}(g_{n})$ via the Law of Large Numbers involving $N$ independent prosumers or via the Central Limit Theorem for independent and dependent random variables [30].
@@ IV-B Market Efficiency with DERA Participation
Consider a transmission network with $I$ buses. At each bus of the transmission network, we assume $M$ PoAs at the distribution network are connected, and $N$ prosumers are aggregated by the proposed DERA model. Denote $U_{in}$ as the concave utility function for the $n$-th prosumer at the $i$-th transmission network bus. ${\bm{g}}_{n}:=(g_{in})_{i\in[I]}$ and ${\bm{d}}_{n}:=(d_{in})_{i\in[I]}$ are respectively the vectors of BTM DG generation and energy consumption for the prosumers. For simplicity, we ignore the number of energy-consuming devices for each prosumer, i.e., $K=1$, in this section. At each transmission bus, we sum up all load-serving entities and generators into one demand function and supply function. The load-serving entity at bus $i$ purchases electricity $e_{i}$ with a concave benefit function $B_{i}(e_{i})$. The generator at bus $i$ produces $p_{i}$ with a convex cost function $C_{i}(p_{i})$. Denote ${\bm{e}}:=(e_{i})_{i\in[I]}$, ${\bm{p}}:=(p_{i})_{i\in[I]}$. ${\bm{f}}\in\mathbb{R}^{L}$ is the line flow limit for $L$ branches of the transmission network. ${\bm{S}}\in\mathbb{R}^{L\times I}$ is the network parameter for DC power flow model.
@@
When prosumers participate in the wholesale market indirectly through the proposed DERA with offer/bid curve (18), social welfare $\mbox{\sf SW}_{\mbox{\rm\tiny DERA}}$ is the optimal value of the convex problem
@@
Proof of this Lemma in our appendix relies on showing that pricing and dispatch results from (19) are at the bidding curve of DERA, i.e., (18). With the optimal dual $\lambda^{\star}\in\mathbb{R}$ for the power balance constraint (19b) and $\hbox{$\mu$}^{\star}\in\mathbb{R}^{L}$ for the line flow limit (19c), the market clearing LMP over $I$ buses is defined by $\hbox{$\pi$}:=\mathbf{1}\lambda^{\star}-{\bm{S}}^{\top}\hbox{$\mu$}^{\star}$, where $\hbox{$\pi$}:=(\pi_{i})_{i\in[I]}$. The prosumer utility in (19a), and constraints for energy consumption and distribution network access in (1)(3) come from DERA’s offer/bid (18).
@@
where ${\cal D}_{in}:=[\max\{\underline{d}_{in},g_{in}-\overline{c}_{in}\},\min\{\bar{d}_{in},g_{in}+\underline{c}_{in}\}]$. The access limits $\overline{c}_{in}$ and $\underline{c}_{in}$ represent the distribution network injection and withdrawal capacities allocated to each prosumer. Detail formulations for $\overline{c}_{in}$ and $\underline{c}_{in}$ are in (56)(57) of our appendix. These values are consistent with (10), ensuring a fair comparison. Under this setup, prosumers participating directly in the wholesale market face the same access constraints at each PoA as those in our proposed DERA model. Solve (21) and obtain the bid/offer curve for prosumer $n$ at bus $i$, i.e.,
@@
The proof is provided in the appendix, which relies on the fact that the proposed DERA has its bidding curve (18) equal to the sum of the prosumer’s bidding curve in (22). From this, we can establish that the wholesale market clearing problem with the direct participation of all prosumers has the same market-clearing results as (19).
@@ V DERA-DSO coordination
All generation and consumption resources aggregated by DERA need to bypass the distribution network to participate in the wholesale market. The DERA aggregation presented in this work ensures that the aggregated DER at each distribution network PoA is bounded by access limits imposed through the distribution network access limit auction in [19]. A DERA submits a bid curve in this auction representing its willingness to acquire access at PoAs. We assume that a DERA is a price taker in the access limit auction. Therefore, the bid-in demand curve for network access from the DERA at a particular PoA is the marginal benefit (profit) from having a DER aggregation under the PoA. The maximum expected profit of DERA is
@@ V-A DERA Benefit Function for Distribution Network Access
The proof is provided in appendix with ${\bm{h}}_{n}(x):=\sum_{k=1}^{K}h_{nk}(x)$, $U_{n}({\bm{h}}_{n}(x))=\sum_{k=1}^{K}U_{nk}(h_{nk}(x))$ from the additive property of the utility function, and
The optimal DERA surplus is decomposed into three terms: one dependent on the withdrawal access $\underline{\phi}_{m}({\underline{C}_{m}})$, one dependent on the injection access $\overline{\phi}_{m}({\overline{C}})$, and one independent of network access. $\varphi({\underline{{\bm{C}}}},{\overline{{\bm{C}}}})$ is separable across injection and withdrawal access over $M$ PoAs. Therefore, DERA can bid separately for the distribution network accesses at different PoAs when coordinating with DSO. At PoA $m$ with less BTM DG, i.e., $\sum_{n=1}^{{\cal N}_{m}}g_{n}\leq\underline{q}_{m}$, DERA’s benefit depends on the withdrawal access. Conversely, if there is more BTM DG, i.e., $\overline{q}_{m}\leq\sum_{n=1}^{{\cal N}_{m}}g_{n}$, DERA’s benefit depends on the injection access. Related simulations are in Sec. VI-D.
@@ V-B Long-Run Equilibrium for Competitive DERA
In a long-run competitive industry, we explore how many DERA can survive. DERAs compete to attract customers, attain distribution network access, and participate in the wholesale market. We assume all DERAs adopt the competitive DER aggregation method in (6). The condition for a competitive long-run equilibrium [31, P193] has two components: (i) the marginal benefit of DERA equals the marginal cost of DSO for providing the distribution network access, and (ii) all DERAs have profits equal to zero, i.e., DERA’s profit in conducting aggregation equals DERA’s payment to acquire distribution network access. Related derivations and simulations are in Sec. VI-E and the appendix.
@@ VI-A Parameter Settings
where $\alpha=\mathdollar 0.4/\mbox{kWh},\beta=\mathdollar 0.1/(\mbox{kWh})^{2}$ [5]. Let the marginal utility $V^{-1}\in[\underline{d},\bar{d}]$ for the consumption limits. [^5] We simulated the case with prosumer device number $K=1$ for simplicity.
We used NEMa and NEMp to represent the DER aggregation under NEM when prosumers were active and passive, respectively. Passive customers are not responsive to the retail prices, but active customers will optimize their energy consumption given the retail price and the BTM DG generations. Based on PG&E residential rate, we set $\pi^{+}=\mathdollar 0.3/\mbox{kWh}$ for the NEM. We assumed $\pi^{-}=\mathbb{E}[\pi]$ and the fixed cost of NEM was covered by extracting fixed payment from DERA, so we simulated with $\pi^{0}=\mathdollar 0$. Gao-Alshehri-Birge (GAB) represented the two-part pricing in [10], which allowed customers to sell BTM DG to the DERA while purchasing energy from its incumbent utility company. Detailed models for NEMa, NEMp, and GAB are provided in Appendix VIII-B. Our DER aggregation method was simulated in Co.NEMa and Co.GAB, competitive with NEMa and GAB, respectively. For Co.NEMa, we set profit ratio $\zeta$ at the upper bound in the proof of Proposition 2. For Co.GAB, we set $\zeta=1.05$ to provide 5% more customer surplus than the GAB competitor.
@@ VI-D Benefit Function of DERA for Distribution Network Access
Figure 5: DERA Benefit function $\varphi$. (Left: withdrawal access $\underline{C}$; right: injection access $\overline{C}$.)
@@ VI-E Long-Run Competitive Equilibrium of DERAs
In the long-run competitive equilibrium analysis with multi-interval aggregation of DERs, we assumed 200 DERAs initially existed and computed the expected number of surviving DERAs in the long run. For simplicity, we assumed DERAs were homogeneous and had the same setting as Sec. VI-D. Prosumers had the same expected DG generation created from the 24-hour roof-top solar data in Pecan Street [33]. [^7] Detail DG trajectories and the long-run equilibrium results for single-interval aggregation are shown in Appendix VIII-G, providing intuitions about long-run equilibrium for multi-interval aggregation here. We multiplied the mean of 24-hour DG by $\epsilon_{1}\in\mathbb{R}_{+}$ to simulate different DG installation capacities and sampled 10,000 random DG scenarios. DERA submitted the benefit function, as in Fig. 5, to acquire hourly distribution network access. Same as [19], the DSO cost function for providing distribution network access was assumed to be the sum of quadratics, $J(x)=\frac{1}{2}bx^{2}+ax$ with $a=\mathdollar 0.009/\mbox{kWh},b=\mathdollar 0.0005/(\mbox{kWh})^{2}$ for both the injection and withdrawal access. We multipied DSO’s cost $J$ by $\epsilon_{2}\in\mathbb{R}_{+}$ to simulate different levels of DSO’s costs.
@@ VII Conclusions
A major challenge to realizing the direct wholesale market participation of DERA is enabling profit-maximizing DERAs to effectively compete with the retail programs offered by utilities in the distribution networks. To this end, this paper considers the competitive DER aggregation of a profit-seeking DERA in the wholesale electricity market. As a wholesale market participant, DERA can both inject and withdraw power from the wholesale market. It is shown that the proposed DERA model maximizes its profit while providing competitive services to its customers with higher surpluses than those offered by the distribution utilities. We also establish that the resulting social welfare from DERA’s participation on behalf of its prosumers is the same as that gained by the direct participation of price-taking prosumers, making the proposed DERA aggregation model optimal in achieving wholesale market efficiency. Additionally, we derive two significant optimal price-quantity bids of DERA, of which one is submitted to the wholesale market, and the other to the distribution network access allocation [19].
An open issue of the proposed aggregation solution is that the payment functions for prosumers are nonlinear and non-uniform. Although each customer is guaranteed to be better off than the competing scheme, two customers producing the same amount may be paid and compensated differently. In other words, the total charge/credits depend not only on the quantity but also on the flexibility of the demand and constraints imposed by the prosumer. Note that a profit-seeking DERA participating in the wholesale electricity market is not subject to the same regulations as a regulated utility. Such non-uniform pricing may be acceptable and has also been proposed in the form of non-uniform fixed charges [10, 7].